# Deep Learning for Protein-Ligand Docking: Are We There Yet?

### v1

@@ Appendix A Broader Impacts
Appendix B Compute Resources
@@
Appendix C Additional Results
@@
In this section, we provide additional results for each baseline method using the Astex Diverse and PoseBusters Benchmark datasets as well as the CASP15 ligand targets. Note that for all violin plots listed in this section, we curate them using combined results across each method’s 3 independent runs (where applicable), in contrast to this section’s bar charts where we instead report mean and standard deviation values across each method’s 3 independent runs.
C.1 Expanded Astex & PoseBusters results
In Figures 7 and 8, we report the ligand RMSD values of each baseline method across the Astex Diverse and PoseBusters Benchmark datasets, with relaxation being applied in the context of the PoseBusters Benchmark dataset. In short, we see that most methods are relatively similar in terms of their ligand RMSD distributions, with RoseTTAFold-All-Atom and our ensembling consensus baseline (i.e., Ensemble (Con)), however, offering more condensed distributions overall. Interestingly, for Astex Diverse, TULIP also appears to produce a uniquely confined ligand RMSD distribution.
@@ C.1 Expanded Astex & PoseBusters results
C.2 Expanded CASP15 results
C.2.1 Overview of expanded results
@@
C.2.2 Multi-ligand RMSD and lDDT-PLI
To start, Figures 9 and 10 report each method’s multi-ligand RMSD and lDDT-PLI distributions with and without relaxation. We see that NeuralPLexer and Ensemble (Con) produce the most tightly bound and accurate RMSD and lDDT-PLI distributions overall.
C.2.3 All single-ligand results
Next, Figures 11, 12, 13, and 14 display each method’s single-ligand CASP15 docking success rates, PoseBusters validity rates, docking RMSD, and docking lDDT-PLI distributions, respectively. In summary, we can make a few respective observations from these plots. (1) DiffDock-L and Ensemble (Con) are the only methods capable of successfully docking any single-ligand CASP15 complexes. (2) AutoDock Vina produces the most PB-valid single-ligand complexes overall, with DiffDock-L shortly behind. (3) DiffDock-L, AutoDock Vina, and Ensemble (Con) appear to achieve the most tightly bound and accurate RMSD distributions. (4) In contrast to (3), only Ensemble (Con) appears to achieve top results in terms of lDDT-PLI compared to the other baseline methods.
C.2.4 Single and multi-ligand results for public targets
Lastly, for completeness and reproducibility, Figures 15, 16, 17, and 18 present corresponding multi-ligand results for the public CASP15 targets, whereas Figures 19, 20, 21, and 22 report corresponding single-ligand results for the public CASP15 targets. Overall, we observe marginal differences between the full and public CASP15 target results for multi-ligand complexes. However, we notice more striking performance drops between the full and public single-ligand CASP15 target results, suggesting that some of the private single-ligand complexes are easier prediction targets than most of the publicly available single-ligand complexes.

### v2

@@ Appendix B Broader impacts
Appendix C Compute resources
@@
Appendix D Documentation for datasets
Below, we provide detailed documentation for each dataset included in our benchmark, summarised in Table 1. Each dataset is freely available for download from the benchmark’s accompanying Zenodo data record [Morehead et al., 2024] under a CC-BY 4.0 license. In lieu of being able to create associated metadata for each of our macromolecular datasets using an ML-focused library such as Croissant [Akhtar et al., 2024] (due to file type compatibility issues), instead, we report structured metadata for our preprocessed datasets using Zenodo’s web user interface [Morehead et al., 2024]. Note that, for all datasets, we authors bear all responsibility in case of any violation of rights regarding the usage of such datasets.
D.1 Astex Diverse Set - Single-Ligand Docking (Difficulty: Easy)
A common drug discovery task is to screen several novel drug-like molecules against a target protein in rapid succession. The Astex Diverse dataset was originally developed with this application in mind, as it features many therapeutically relevant 3D molecules for computational modeling.
  • Motivation Several downstream drug discovery efforts rely on having access to high-quality molecular data for docking.
  • Collection For this dataset, which was originally compiled by Hartshorn et al. [2007], we adopt the version further prepared by Buttenschoen et al. [2024].
  • Composition The dataset consists of 85 (80) single-ligand protein complexes (for which we could obtain high-accuracy predicted protein structures using AlphaFold/ESMFold).
  • Hosting Our preprocessed version of the dataset (https://doi.org/10.5281/zenodo.11477766) can be downloaded from the benchmark’s Zenodo data record at https://zenodo.org/records/11477766/files/astex_diverse_set.tar.gz.
  • Licensing We have released our preprocessed version of the dataset under a CC-BY 4.0 license. The original dataset is available under a CC-BY 4.0 license on Zenodo [Buttenschoen et al., 2023].
  • Maintenance We will announce any errata discovered in or changes made to the dataset using the benchmark’s GitHub repository at https://github.com/BioinfoMachineLearning/PoseBench.
  • Uses This dataset of holo (and predicted-apo) protein PDB and holo ligand SDF files can be used for single-ligand docking or protein-ligand structure generation.
  • Metric Ligand RMSD $\leq$ 2 Å & PoseBusters-Valid (PB-Valid).
D.2 PoseBusters Benchmark Set - Single-Ligand Docking (Difficulty: Intermediate)
Like the Astex Diverse dataset, the PoseBusters Benchmark dataset was originally developed for docking individual ligands to target proteins. However, this dataset features a larger and more challenging collection of protein-ligand complexes for computational modeling.
  • Motivation Data sources of challenging single-ligand protein complexes for molecular docking are critical for the development of future docking methods.
  • Collection For this dataset, we adopt the version introduced by Buttenschoen et al. [2024].
  • Composition The dataset consists of 308 (280) single-ligand protein complexes (for which we could obtain high-accuracy predicted protein structures using AlphaFold/ESMFold).
  • Hosting Our preprocessed version of the dataset (https://doi.org/10.5281/zenodo.11477766) can be downloaded from the benchmark’s Zenodo data record at https://zenodo.org/records/11477766/files/posebusters_benchmark_set.tar.gz.
  • Licensing We have released our preprocessed version of the dataset under a CC-BY 4.0 license. The original dataset is available under a CC-BY 4.0 license on Zenodo [Buttenschoen et al., 2023].
  • Maintenance We will announce any errata discovered in or changes made to the dataset using the benchmark’s GitHub repository at https://github.com/BioinfoMachineLearning/PoseBench.
  • Uses This dataset of holo (and predicted-apo) protein PDB and holo ligand SDF files can be used for single-ligand docking or protein-ligand structure generation.
  • Metric Ligand RMSD $\leq$ 2 Å & PoseBusters-Valid (PB-Valid).
D.3 DockGen Set - Single-Ligand Docking (Difficulty: Challenging)
The DockGen dataset was originally developed for docking individual ligands to target proteins in the context of novel protein binding pockets. As such, this dataset is useful for evaluating how well each baseline method can generalize to distinctly different binding pockets compared to those on which it may have been trained.
  • Motivation Data sources of protein-ligand complexes representing novel single-ligand binding pockets are critical for the development of generalizable docking methods.
  • Collection For this dataset, we adopt the version introduced by Corso et al. [2024a].
[…]
  • Hosting Our preprocessed version of the dataset (https://doi.org/10.5281/zenodo.11477766) can be downloaded from the benchmark’s Zenodo data record at https://zenodo.org/records/11477766/files/dockgen_set.tar.gz.
  • Licensing We have released our preprocessed version of the dataset under a CC-BY 4.0 license. The original dataset is available under an MIT license on Zenodo [Corso et al., 2024b].
  • Maintenance We will announce any errata discovered in or changes made to the dataset using the benchmark’s GitHub repository at https://github.com/BioinfoMachineLearning/PoseBench.
  • Uses This dataset of holo (and predicted-apo) protein PDB and holo ligand PDB files can be used for single-ligand docking or protein-ligand structure generation.
  • Metric Ligand RMSD $\leq$ 2 Å & PoseBusters-Valid (PB-Valid).
D.4 CASP15 Set - Multi-Ligand Docking (Difficulty: Challenging)
As the most complex of our benchmark’s four test datasets, the CASP15 protein-ligand interaction dataset was created to represent the new protein-ligand modeling category in the 15th Critical Assessment of Structure Prediction (CASP) competition. Whereas the Astex Diverse and PoseBusters Benchmark datasets feature solely single-ligand protein complexes, the CASP15 dataset provides users with a variety of challenging organic (e.g., drug molecules) and inorganic (e..g., ion) cofactors for multi-ligand biomolecular modeling.
  • Motivation Multi-ligand evaluation datasets for molecular docking provide the rare opportunity to assess how well baseline methods can model intricate protein-ligand interactions while avoiding troublesome protein-ligand and ligand-ligand steric clashes. Additionally, more accurate modeling of multi-ligand complexes in future works may lead to improved techniques for computational enzyme design and regulation [Stärk et al., 2023].
  • Collection For this dataset, we manually collect each publicly and privately available CASP15 protein-bound ligand complex structure compatible with protein-ligand (e.g., non-RNA) benchmarking.
  • Composition The dataset consists of 102 (86) fragment ligands contained within 19 (15) separate (publicly available) protein complexes, of which 6 (2) and 13 (2) of these complexes are single and multi-ligand complexes, respectively.
  • Hosting Our preprocessed version of (the publicly available version of) this dataset (https://doi.org/10.5281/zenodo.11477766) can be downloaded from the benchmark’s Zenodo data record at https://zenodo.org/records/11477766/files/casp15_set.tar.gz.
  • Licensing We have released our preprocessed version of the (public) dataset under a CC-BY 4.0 license. The original (public) dataset is free for download via the RCSB PDB [Bank, 1971].
  • Maintenance We will announce any errata discovered in or changes made to the dataset using the benchmark’s GitHub repository at https://github.com/BioinfoMachineLearning/PoseBench.
  • Uses This dataset of holo (and predicted-apo) protein PDB and holo ligand PDB files can be used for multi-ligand docking or protein-ligand structure generation.
  • Metric (Fragment) Ligand RMSD $\leq$ 2 Å & (Complex) PoseBusters-Valid (PB-Valid).
Figure 7: Astex, PoseBusters, DockGen, and CASP15 dataset analysis of protein-ligand interactions.
Appendix E Analysis of protein-ligand interactions
Inspired by a similar analysis presented in the PoseCheck benchmark [Harris et al., 2023], in this section, we study the frequency of different types of protein-ligand interactions such as Van der Waals contacts and hydrophobic interactions occurring natively within the Astex Diverse, PoseBusters Benchmark, DockGen, and CASP15 datasets, respectively. In particular, these measures allow us to better understand the diversity of interactions each baseline method within the PoseBench benchmark is tasked to model, within the context of each test dataset. Figure 7 displays the results of this analysis.
Overall, we find that the Astex Diverse, PoseBusters Benchmark, and DockGen datasets contain similar types and frequencies of interactions, with the PoseBusters Benchmark and DockGen datasets containing slightly more hydrogen bond acceptors ($\sim$3 vs 1) and Van der Waals contacts ($\sim$13 vs 8) on average compared to the Astex Diverse dataset. However, we note a significant difference in interaction types and frequencies between the CASP15 dataset and the three other datasets. Specifically, we find it contains a significantly higher proportion of hydrogen bond acceptors and donors ($\sim$40), Van der Waals contact ($\sim$200), and hydrophobic interactions ($\sim$15) on average. Particularly interesting to note is the CASP15 dataset’s bimodal distribution of hydrophobic interactions, suggesting that the dataset contains two primary classes of interacting ligands giving rise to hydrophobic interactions. One possible explanation for this phenomenon is that the CASP targets, in contrast to the Astex Diverse, PoseBusters Benchmark, and DockGen targets, consist of a variety of both organic (e.g., drug-like molecules) and inorganic (e.g., metal) cofactors.
Appendix F Additional results
@@
In this section, we provide additional results for each baseline method using the Astex Diverse, PoseBusters Benchmark, and DockGen datasets as well as the CASP15 ligand targets. Note that for all violin plots listed in this section, we curate them using combined results across each method’s three independent runs (where applicable), in contrast to this section’s bar charts where we instead report mean and standard deviation values across each method’s three independent runs.
F.1 DockGen results
DockGen dataset. The DockGen dataset [Corso et al., 2024a] contains 189 diverse single-ligand protein complexes, each representing a novel type of protein-ligand binding pocket. This dataset can be considered the most difficult single-ligand benchmark set since its protein binding sites are distinctly different from those found in the training datasets of most deep learning-based docking methods to date.
[…]
Figure 8: DockGen dataset results for successful single-ligand docking with relaxation.
Figure 9: DockGen dataset results for single-ligand docking RMSD.
Benchmark results. Figures 8 and 9 show that DiffDock-L, NeuralPLexer, and Ensemble (Con) provide the best pocket generalization capabilities compared to all other baseline methods. In particular, Ensemble (Con) seems to provide the most stable performance in this setting. However, we note that none of the baseline methods generate any PB-valid ligand conformations, suggesting that such "correct" poses are approximately accurate yet physically implausible in certain measurable ways.
Figure 10: Pocket-only PoseBusters dataset results for successful single-ligand docking with relaxation.
Figure 11: Pocket-only PoseBusters dataset results for single-ligand docking RMSD.
F.2 Expanded Astex & PoseBusters results
F.2.1 Pocket-only PoseBusters results
Figures 10 and 11 illustrate the impact of reducing the binding pocket search space of each baseline docking method by providing each method with alternative versions of the predicted PoseBusters Benchmark protein structures that have been cropped to contain only ligand-interacting (< 4 Å heavy atom distance) protein residues and their (7) sequence-adjacent neighbors. Overall, we find that performing such pocket-level docking increases the docking success rates and favorably narrows the ligand RMSD distributions of DiffDock-L, NeuralPLexer, RoseTTAFold-All-Atom, and Ensemble (Con), whereas for all other baselines, performance is either maintained or degraded marginally. This finding highlights that methods such as DiffDock-L, NeuralPLexer, and RoseTTAFold-All-Atom are better at identifying the correct local structural conformations of each ligand compared to other baseline methods such as FABind and DynamicBind.
F.2.2 Astex & PoseBusters RMSD results
@@
In Figures 12 and 13, we report the ligand RMSD values of each baseline method across the Astex Diverse and PoseBusters Benchmark datasets, with relaxation being applied in the context of the PoseBusters Benchmark dataset. In short, we see that most methods are relatively similar in terms of their ligand RMSD distributions, with RoseTTAFold-All-Atom and our ensembling consensus baseline (i.e., Ensemble (Con)), however, offering more condensed distributions overall. Interestingly, for Astex Diverse, TULIP also appears to produce a uniquely confined ligand RMSD distribution.
F.3 Expanded CASP15 results
F.3.1 Overview of expanded results
@@
F.3.2 Multi-ligand RMSD and lDDT-PLI
To start, Figures 14 and 15 report each method’s multi-ligand RMSD and lDDT-PLI distributions with and without relaxation. We see that NeuralPLexer and Ensemble (Con) produce the most tightly bound and accurate RMSD and lDDT-PLI distributions overall.
F.3.3 All single-ligand results
Next, Figures 16, 17, 18, and 19 display each method’s single-ligand CASP15 docking success rates, PoseBusters validity rates, docking RMSD, and docking lDDT-PLI distributions, respectively. In summary, we can make a few respective observations from these plots. (1) DiffDock-L and Ensemble (Con) are the only methods capable of successfully docking any single-ligand CASP15 complexes. (2) AutoDock Vina produces the most PB-valid single-ligand complexes overall, with DiffDock-L shortly behind. (3) DiffDock-L, AutoDock Vina, and Ensemble (Con) appear to achieve the most tightly bound and accurate RMSD distributions. (4) In contrast to (3), only Ensemble (Con) appears to achieve top results in terms of lDDT-PLI compared to the other baseline methods.
F.3.4 Single and multi-ligand results for public targets
Lastly, for completeness and reproducibility, Figures 20, 21, 22, and 23 present corresponding multi-ligand results for the public CASP15 targets, whereas Figures 24, 25, 26, and 27 report corresponding single-ligand results for the public CASP15 targets. Overall, we observe marginal differences between the full and public CASP15 target results for multi-ligand complexes. However, we notice more striking performance drops between the full and public single-ligand CASP15 target results, suggesting that some of the private single-ligand complexes are easier prediction targets than most of the publicly available single-ligand complexes.